On July 1, 2013, Watson Co. paid $215,000 for all of the stock of Squire, Inc. On that date, book values of Squire's assets and liabilities were $200,000 and $45,000, respectively. The fair values of the assets and liabilities were $210,000 and $35,000, respectively. What is the amount of goodwill at date of acquisition?
Answer #1

Fair value of the assets - Fair value of the liabilities = 210,000 - 35,000 = 175,000

Good will on acquisition = Price paid - fair value = 215,000 - 175,000 = 40,000